OpenStudy (anonymous):

perform the requested operation or operations. f(x) = x minus five divided by eight.; g(x) = 8x + 5, find g(f(x))

OpenStudy (debbieg):

As written, your f(x) is \[f(x)=x-\frac{ 5 }{8}\], but I suspect that you actually want \[f(x)=\frac{ x-5 }{ 8 }\]?? In either case, all you're doing here is a composition of functions. g(f(x)) means to use the rule established by the function g and apply it to the function f(x), that is, evaluate g(x) using f(x) as your "input" - the independent variable x. So substitute the expression\[\frac{ x-5 }{ 8 }\] into g(x) for the variable x, and then simplify. Make sense? Let me know what you get. :)